# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v2

@@ B.2 Additional Quantitative Results
B.3 Additional Qualitative Results
To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems. For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels. For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair. To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates. The following figures (Figure 6, Figure 7, Figure 8, Figure 9) illustrate that DPBridge achieves visually satisfactory results across different tasks. Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape (Cordts et al., 2016) dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ (Lee et al., 2020) dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel (Butler et al., 2012) dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 9: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics (Sxela, 2021) dataset.

### v3

@@ E.1 Additional Quantitative Results
E.2 Additional Qualitative Results
To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, and style transfer. Qualitative results are shown in Figure 7, Figure 8, Figure 9, and Figure 10. These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem, and replacing the paired ground truth data accordingly. For semantic segmentation, we convert discrete labels into color maps and train the bridge model to predict colors in RGB space rather than class indices. For optical flow, we define the bridge process between the flow image and the first frame of the input RGB pair. To accommodate this, we modify the UNet architecture to condition on both RGB frames during inference and represent flow as a color-encoded image, enabling prediction in RGB space rather than raw uv coordinates. Across all tasks, DPBridge produces visually compelling results, demonstrating its versatility as a unified framework that can be easily adapted to a range of dense prediction scenarios. We note that these results are qualitative; quantitative evaluation and task-specific architectural refinements are left for future work.
Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [12] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 8: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [30] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 9: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [5] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [48] dataset.
Appendix F Broader Impacts
DPBridge aims to advance the field of dense prediction through a generative modeling perspective, integrating the strengths of diffusion bridge formulation with visual priors from pretrained backbones. Our proposed method is a unified framework that can be easily extend to different tasks, including depth prediction, surface normal estimation, semantic segmentation, optical flow, etc., which will benefit the whole dense prediction research community. In addition, improved generalization and precision in dense prediction can also benefit a wide range of real-world applications, such as autonomous driving, augmented reality, and digital content creation.